how many moles of o2 are produced when 1.26 moles of h2o is reacted

1
Expert's answer
2022-05-13T08:56:04-0400

2H2O = O2 + 2H2 (electrolysis)

According to the equation, n (O2) = 1/2 x n (H2O) = 1/2 x 1.26 = 0.63 mol